OpenStudy (anonymous):

Find |AB| if A=(1,-4) and B=(-2,7). Also, how would I find the direction of AB?

OpenStudy (anonymous):

|AB|=sqrt ((1+2)^2+(-4-7)^2) = sqrt 130

OpenStudy (anonymous):

Thank you! Also do you know how to find the direction?

OpenStudy (anonymous):

also vector AB = vector B - vector A= (-2-1)i +(7+4)j=-3i + 11j so unit vector along AB =vector AB /|AB| =-3i/sqrt 130 + 11j/ sqrt 130

OpenStudy (anonymous):

Isn't AB (1--2),(-4-7), so it would be <3,-11> or 3i-11j?

OpenStudy (amistre64):

if A and B are points; then B-A is the direction of the vector FROM A to B

OpenStudy (amistre64):

(..) tends to mean points <...> tends to indicate vecotrs but thats not a universal notation

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

youre welcome